OpenStudy (mvpriest6):

when Y=3, x=9, and x varies directly with y, match each value of y with the corresponding value of x terms: y=5, y=4, y=2, y=6 definitions: x=12, x=18, x=15, x=6 please help

OpenStudy (freckles):

well first translate "x varies directly with y" into mathematical equation

OpenStudy (mvpriest6):

so y=x

OpenStudy (freckles):

there should be a constant somewhere in there

OpenStudy (mvpriest6):

y=x+b?

OpenStudy (freckles):

a constant multiple

OpenStudy (freckles):

"x varies directly with y" translates to ky=x

OpenStudy (mvpriest6):

oh ok now what

OpenStudy (freckles):

find your constant k first

OpenStudy (freckles):

you are given a point on ky=x

OpenStudy (freckles):

when Y=3, x=9, and x varies directly with y the point being (x=9,y=3) plug that into ky=x to find k

OpenStudy (mvpriest6):

k3=9

OpenStudy (freckles):

solve for k

OpenStudy (mvpriest6):

3+3=9 so k=3?

OpenStudy (freckles):

well 3+3 doesn't equal 9 but 3(3)=9 so k=3 since k=9/3

OpenStudy (freckles):

\[3y=x\]

OpenStudy (mvpriest6):

yea sorry so how do I figure out the answers

OpenStudy (freckles):

now plug in your y's to match the x's you are given to them

OpenStudy (mvpriest6):

I see if I can multiply the 3 by the x's to equal y? so 3*5

OpenStudy (freckles):

3y=x you are replace y with y not y with x you are multiply the y's by 3

OpenStudy (mvpriest6):

so the other way around

OpenStudy (freckles):

terms: y=5, y=4, y=2, y=6 take all these things called y and multiply it by 3 just like the equation 3y=x says to

OpenStudy (mvpriest6):

oh ok thx I understand now